A 250g copper pipe is heated from 283k to 304k what is the energy needed to heat the pipe?the specific heat capacity of copper is 390J/kg-1 degree Celsius-1
1
Expert's answer
2020-11-23T10:26:25-0500

We can find the energy needed to heat the copper pipe from the formula:


"Q=mc\Delta T,""Q=0.25\ kg\cdot 390\ \dfrac{J}{kg\cdot K}\cdot(304\ K-283\ K)=2047.5\ J."

Answer:

"Q=2047.5\ J."